OpenStudy (anonymous):

Simplify 12x^3/16-x^2 _______________ (divided by) 3x^4/20+5x

OpenStudy (anonymous):

\[12x^3 \div 16-x^2 /3x^4 \div20+5x\]

OpenStudy (mathstudent55):

Is this it? \(\huge \dfrac{\frac{12x^3}{16} - x^2}{\frac{3x^4}{20}+ 5x} \)

OpenStudy (mathstudent55):

If so, then first, reduce the fraction in the numerator. You will have a new smaller denominator than 16. Then what is the LCM of the new denominator that is less than 16 and 20? Multiply the main fraction by that denominator over itself to eliminate the small fractions.

OpenStudy (mathstudent55):

Then factor the expressions int eh numerator and denominator.

OpenStudy (triciaal):

Please clarify the original.